Calculate the potential energy of a 4.5 kg bowling ball that is on the ground. Assume that g = 9.8 m/s2


1
Expert's answer
2021-01-15T13:46:34-0500

By the definition of the potential energy, we have:


"PE=mgh=4.5\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot 0\ m=0."

Answer:

"PE=0."